# Decentralized Nonconvex Optimization under Heavy-Tailed Noise: Normalization and Optimal Convergence

### v1

@@ B.2 Intermediate Lemmas
$\mathbb{E}\big{[}\|\sum_{k=0}^{t}\beta^{t-k}(1-\beta)\tilde{\bm{s}}^{k}\|\mid\mathcal{F}_{t-1}]$
=\ = $\mathbb{E}\big{[}\|\sum_{k=0}^{t}\beta^{t-k}(1-\beta)(\bm{g}(\bm{x}^{t},\bm{\xi}^{t,b})-\nabla F(\bm{x}^{t}))\|\mid\mathcal{F}_{t-1}\big{]}$
[…]
$\overset{(i)}{\leq}$ $2\sqrt{2}\Big{(}\mathbb{E}\big{[}\sum_{k=0}^{t}\beta^{(t-k)p}(1-\beta)^{p}\|\bm{g}(\bm{x}^{t},\bm{\xi}^{t})-\nabla F(\bm{x}^{t})\|^{p}\big{]}\mid\mathcal{F}_{t-1}\Big{)}^{\frac{1}{p}}$
$\overset{(ii)}{\leq}$ $2\sqrt{2}\Big{(}\mathbb{E}\big{[}\sum_{k=0}^{t}\beta^{(t-k)p}(1-\beta)^{p}\big{(}\sum_{i=1}^{n}\|\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t})\|\big{)}^{p}\big{]}\mid\mathcal{F}_{t-1}\Big{)}^{\frac{1}{p}}$
$\overset{(iii)}{\leq}$ $2\sqrt{2}\Big{(}\mathbb{E}\big{[}\sum_{k=0}^{t}\sum_{i=1}^{n}\beta^{(t-k)p}(1-\beta)^{p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t})\|^{p}\big{]}\mid\mathcal{F}_{t-1}\Big{)}^{\frac{1}{p}}$
=\ = $2\sqrt{2}\Big{(}\mathbb{E}\big{[}\sum_{i=1}^{n}(1-\beta)^{p}n^{p-1}\|\nabla\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t})\|^{p}\mid\mathcal{F}_{t-1}\big{]}$
$\quad+\sum_{k=0}^{t-1}\sum_{i=1}^{n}\beta^{(t-k)p}(1-\beta)^{p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t})\|^{p}\Big{)}^{\frac{1}{p}}$
[…]
where in $(i)$ we used Jensen’s inequality, and in $(ii),(iii)$ we used relations $4$,and $5$ in Lemma 1, respectively. Applying the above arguments from $\mathcal{F}_{t-2}$ to $\mathcal{F}_{0}$, we obtain
@@
$\overset{(i)}{\leq}$ $n\sum_{k=0}^{t}\beta^{t-k+1}\Big{(}\frac{2\alpha\lambda}{1-\lambda}+\alpha\Big{)}L.$
@@
Now we are ready to proof our main theorems.
@@
$\overset{(i)}{\leq}$ $f(\bar{\bm{x}}^{0})-f_{*}+2\|\nabla f(\bar{\bm{x}}^{0})\|\cdot\frac{\alpha}{1-\beta}+\frac{4\sqrt{2}\sigma}{n^{1-\frac{1}{p}}}\cdot\alpha(1-\beta)^{1-\frac{1}{p}}T+\frac{4L}{1-\lambda}\cdot\frac{\alpha^{2}T}{1-\beta}+\frac{2L}{1-\lambda}\cdot\alpha^{2}T$
@@
$\begin{split}&\frac{1}{nT}\sum_{t=0}^{T-1}\sum_{i=1}^{n}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{i}^{t})\|]\\ \leq\ &\frac{\Delta_{0}}{\alpha T}+\frac{2\|\nabla f(\bar{\bm{x}}^{0})\|}{(1-\beta)T}+4\sqrt{2}\sigma\cdot\frac{(1-\beta)^{1-\frac{1}{p}}}{n^{1-\frac{1}{p}}}+\frac{4L}{1-\lambda}\cdot\frac{\alpha}{1-\beta}+\Big{(}\frac{1+2L}{1-\lambda}+\frac{L}{2}\Big{)}\alpha\\ &\quad+\frac{2\sqrt{2}\sigma}{(1-\lambda)^{\frac{1}{p}}}\cdot n^{\frac{1}{2}}\big{(}\frac{1}{\beta}-1\big{)}+\frac{\|\nabla F(\mathbf{1}_{n}\otimes\bar{\bm{x}}^{0})\|}{1-\lambda}\cdot\frac{\frac{1}{\beta}-1}{n^{\frac{1}{2}}}+\frac{2\sqrt{2}\sigma}{1-\lambda}\cdot n^{\frac{1}{2}}\big{(}\frac{1}{\beta}-1\big{)}(1-\beta)^{1-\frac{1}{p}}\\ &\quad+\frac{2L}{(1-\lambda)^{2}}\cdot n^{\frac{1}{2}}\alpha\\ \overset{(i)}{\leq}\ &\frac{\Delta_{0}}{\alpha T}+\frac{2\|\nabla f(\bar{\bm{x}}^{0})\|}{(1-\beta)T}+4\sqrt{2}\sigma\cdot\frac{(1-\beta)^{1-\frac{1}{p}}}{n^{1-\frac{1}{p}}}+\frac{4L}{1-\lambda}\cdot\frac{\alpha}{1-\beta}+\frac{1+2.5L}{1-\lambda}\alpha\\ &\quad+\frac{20\sqrt{2}\sigma}{(1-\lambda)^{\frac{1}{p}}}\cdot n^{\frac{1}{2}}(1-\beta)+\frac{10\|\nabla F(\mathbf{1}_{n}\otimes\bar{\bm{x}}^{0})\|}{1-\lambda}\cdot\frac{1-\beta}{n^{\frac{1}{2}}}+\frac{20\sqrt{2}\sigma}{1-\lambda}\cdot n^{\frac{1}{2}}(1-\beta)^{2-\frac{1}{p}}+\frac{2L}{(1-\lambda)^{2}}\cdot n^{\frac{1}{2}}\alpha\\ \overset{(ii)}{\leq}\ &O\Big{(}\frac{\Delta_{0}}{T}+\frac{2\|\nabla f(\bar{\bm{x}}^{0})\|}{(1-\beta)T}+4\sqrt{2}\sigma\cdot\frac{(1-\beta)^{1-\frac{1}{p}}}{n^{1-\frac{1}{p}}}+\sqrt{\frac{L\Delta_{0}}{(1-\lambda)(1-\beta)T}}+\sqrt{\frac{(1+2.5L)\Delta_{0}}{(1-\lambda)T}}\\ &\quad+\frac{20\sqrt{2}\sigma}{(1-\lambda)^{\frac{1}{p}}}\cdot n^{\frac{1}{2}}(1-\beta)+\frac{10\|\nabla F(\mathbf{1}_{n}\otimes\bar{\bm{x}}^{0})\|}{1-\lambda}\cdot\frac{1-\beta}{n^{\frac{1}{2}}}+\frac{\sigma}{1-\lambda}\cdot n^{\frac{1}{2}}(1-\beta)^{2-\frac{1}{p}}+\sqrt{\frac{n^{\frac{1}{2}}L\Delta_{0}}{(1-\lambda)^{2}T}}\Big{)}\\ \overset{(iii)}{\leq}\ &O\Big{(}\frac{\Delta_{0}}{T}+\frac{\|\nabla f(\bar{\bm{x}}^{0})\|}{T^{\frac{2p-2}{3p-2}}}+\frac{\sigma}{n^{1-\frac{1}{p}}T^{\frac{p-1}{3p-2}}}+\sqrt{\frac{L\Delta_{0}}{(1-\lambda)T^{\frac{2p-2}{3p-2}}}}+\sqrt{\frac{(1+2.5L)\Delta_{0}}{(1-\lambda)T}}\\ &\quad+\frac{\sigma n^{\frac{1}{2}}}{(1-\lambda)^{\frac{1}{p}}T^{\frac{p}{3p-2}}}+\frac{\|\nabla F(\mathbf{1}_{n}\otimes\bar{\bm{x}}^{0})\|}{(1-\lambda)n^{\frac{1}{2}}T^{\frac{p}{3p-2}}}+\frac{\sigma}{1-\lambda}\frac{n^{\frac{1}{2}}}{T^{\frac{2p-1}{3p-2}}}+\sqrt{\frac{n^{\frac{1}{2}}L\Delta_{0}}{(1-\lambda)^{2}T}}\Big{)}\\ \end{split}$ (29)
@@
$\alpha=\min\Big{(}1,\ \max\Big{(}\sqrt{\frac{\Delta_{0}(1-\beta)(1-\lambda)}{4LT}},\sqrt{\frac{\Delta_{0}(1-\lambda)}{(1+2.5L)T}},\sqrt{\frac{(1-\lambda)^{2}\Delta_{0}}{2n^{\frac{1}{2}}LT}}\Big{)}\Big{)},$ (30)
@@
Rearranging above terms lead to the desired upper bound. ∎
@@ C.2 Additional details for synthetic experiments
$\ell(r)=\begin{cases}\frac{c^{2}}{6}\left(1-\left[1-\left(\frac{r}{c}\right)^{2}\right]^{3}\right)&\text{ if }|r|\leq c,\\ \frac{c^{2}}{6}&\text{ otherwise }\end{cases},$
@@ C.3 Additional details for decentralized training of Transformers
Transformer architecture. We consider the following decoder-only Transformer model (GPT): vocabulary size is 10208, context length is 64, embedding size is 128, number of attention heads is 4, number of attention layers is 2, the linear projection dimension within attention block is 512, and LayerNorm is applied after the 2nd attention block. The total number of parameter of this model is 3018240.
Training losses. In Figure 5, we also plot the training losses of all algorithms in our decentralized training experiment. Consistent with the validation losses, GT-NSGDm is the only algorithm that achieves convergence to the optimum, with final average losses of 0.100, 0.096, and 0.091 on ring, directed exponential, and complete graphs, respectively.
(a) Ring Graph
(b) Directed Exponential Graph
(c) Complete Graph
Figure 5: Comparison of graph-wide average training losses in decentralized training of Transformer models, over ring, directed exponential, and complete graphs.
@@
Hardware. We simulate the distributed training on one NVIDA H100 GPU, using PyTorch 3.2 with CUDA 12. The total hyperparameter search and training procedure took around 100 GPU hours.

### v2

@@ B.2 Intermediate Lemmas
$\begin{split}&\mathbb{E}\big[\|\sum_{k=0}^{t}\beta^{t-k}(1-\beta)\tilde{\bm{s}}^{k}\|]\\ =\ &\mathbb{E}\big[\|\sum_{k=0}^{t}\beta^{t-k}(1-\beta)(\bm{g}(\bm{x}^{t},\bm{\xi}^{t,b})-\nabla F(\bm{x}^{t}))\|\big]\\ \leq\ &2\sqrt{2}\mathbb{E}\big[\big(\sum_{k=0}^{t}\beta^{(t-k)p}(1-\beta)^{p}\|\bm{g}(\bm{x}^{t},\bm{\xi}^{t})-\nabla F(\bm{x}^{t})\|^{p}\big)^{\frac{1}{p}}\big].\end{split}$ (32)
In addition,
$\begin{split}&2\sqrt{2}\mathbb{E}\big[\big(\sum_{k=0}^{t}\beta^{(t-k)p}(1-\beta)^{p}\|\bm{g}(\bm{x}^{t},\bm{\xi}^{t})-\nabla F(\bm{x}^{t})\|^{p}\big)^{\frac{1}{p}}\mid\mathcal{F}_{t-1}\big]\\ \overset{(i)}{\leq}\ &2\sqrt{2}\Big(\mathbb{E}\big[\sum_{k=0}^{t}\beta^{(t-k)p}(1-\beta)^{p}\|\bm{g}(\bm{x}^{t},\bm{\xi}^{t})-\nabla F(\bm{x}^{t})\|^{p}\big]\mid\mathcal{F}_{t-1}\Big)^{\frac{1}{p}}\\ \overset{(ii)}{\leq}\ &2\sqrt{2}\Big(\mathbb{E}\big[\sum_{k=0}^{t}\beta^{(t-k)p}(1-\beta)^{p}\big(\sum_{i=1}^{n}\|\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t})\|\big)^{p}\big]\mid\mathcal{F}_{t-1}\Big)^{\frac{1}{p}}\\ \overset{(iii)}{\leq}\ &2\sqrt{2}\Big(\mathbb{E}\big[\sum_{k=0}^{t}\sum_{i=1}^{n}\beta^{(t-k)p}(1-\beta)^{p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t})\|^{p}\big]\mid\mathcal{F}_{t-1}\Big)^{\frac{1}{p}}\\ =\ &2\sqrt{2}\Big(\mathbb{E}\big[\sum_{i=1}^{n}(1-\beta)^{p}n^{p-1}\|\nabla\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t})\|^{p}\mid\mathcal{F}_{t-1}\big]\\ &\quad+\sum_{k=0}^{t-1}\sum_{i=1}^{n}\beta^{(t-k)p}(1-\beta)^{p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t})\|^{p}\Big)^{\frac{1}{p}}\\ \leq\ &2\sqrt{2}\Big((1-\beta)^{p}n^{p}\sigma^{p}+\sum_{k=0}^{t-1}\sum_{i=1}^{n}\beta^{(t-k)p}(1-\beta)^{p}n^{p-1}\|\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x}_{i}^{t})\|^{p}\Big)^{\frac{1}{p}},\end{split}$ (33)
where in $(i)$ we used Jensen’s inequality, and in $(ii),(iii)$ we used relations $4$,and 55 in Lemma 1, respectively. Based on (32), taking expectations on both sides of (33), and applying the above arguments from $\mathcal{F}_{t-2}$ to $\mathcal{F}_{0}$, we obtain
@@
$\overset{(i)}{\leq}\$$n\sum_{k=0}^{t}\beta^{t-k+1}\Big(\frac{2\alpha\lambda}{1-\lambda}+\alpha\Big)L.$
@@
Now we are ready to prove our main theorems.
@@
$\overset{(i)}{\leq}\$$f(\bar{\bm{x}}^{0})-f_{*}+2\|\nabla f(\bar{\bm{x}}^{0})\|\cdot\frac{\alpha}{1-\beta}+\frac{4\sqrt{2}\sigma}{n^{1-\frac{1}{p}}}\cdot\alpha(1-\beta)^{1-\frac{1}{p}}T+\frac{4L}{1-\lambda}\cdot\frac{\alpha^{2}T}{1-\beta}+\frac{2L}{1-\lambda}\cdot\alpha^{2}T$
@@
[…]
@@
$\alpha=\min\Big(1,\ \sqrt{\frac{\Delta_{0}(1-\beta)(1-\lambda)}{4LT}},\sqrt{\frac{\Delta_{0}(1-\lambda)}{3.5LT}},\sqrt{\frac{(1-\lambda)^{2}\Delta_{0}}{2n^{\frac{1}{2}}LT}}\Big),$ (36)
@@
Rearranging above terms leads to the desired upper bound. ∎
@@ C.1 Baseline descriptions
DSGD-GClip $\bm{x}_{i}^{t+1}=\sum_{r=1}^{n}w_{ir}\bm{x}_{r}^{t}-\alpha\operatorname{\textsf{clip}}(g_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t}),\tau)$ $\alpha,\tau$: stepsize $\alpha$, and $\ell_{2}$ clipping levels $\tau$
DSGD-CClip $\bm{x}_{i}^{t+1}=\sum_{r=1}^{n}w_{ir}\bm{x}_{r}^{t}-\alpha\operatorname{\textsf{clip}}(g_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t}),\tau)$ $\alpha,\tau$: stepsize $\alpha$, and component-wise clipping levels $\tau$
@@
GT-Adam $\begin{array}[]{l}\bm{m}_{i}^{t+1}=\beta_{1}\bm{m}_{i}^{t}+(1-\beta_{1})\bm{s}_{i}^{t}\\ \bm{v}_{i}^{t+1}=\min\big(\beta_{2}\bm{v}_{i}^{t}+(1-\beta_{2})\bm{s}_{i}^{t}\odot\bm{s}_{i}^{t},G\big)\\ \bm{x}_{i}^{t+1}=\sum_{r=1}^{n}w_{ir}\bm{x}_{r}^{t}-\alpha\frac{\bm{m}_{i}^{t+1}}{\sqrt{\bm{v}_{i}^{t+1}+\epsilon}}\\ \bm{g}_{i}^{t+1}=\nabla f_{i}(\bm{x}_{i}^{t+1})\\ \bm{s}_{i}^{t+1}=\sum_{r=1}^{n}w_{ir}\bm{s}_{r}^{t}+\bm{g}_{i}^{t+1}-\bm{g}_{i}^{t}\end{array}$ $\alpha,G$: constant stepsize $\alpha$, and upper bound $G$, stabilization factor $\epsilon$
QG-DSGDm $\begin{array}[]{l}\bm{m}_{i}^{t+1}=\beta\hat{\bm{m}}_{i}^{t}+g_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})\\ \bm{x}_{i}^{t+1}=\sum_{r=1}^{n}w_{ir}\big(\bm{x}_{i}^{t}-\eta\bm{m}_{i}^{t+1}\big)\\ \bm{d}_{i}^{t}=(\bm{x}_{i}^{t+1}-\bm{x}_{i}^{t})/\eta\\ \hat{\bm{m}}_{i}^{t+1}=\mu\hat{\bm{m}}_{i}^{t}+(1-\mu)\bm{d}_{i}^{t}\end{array}$ $\eta,\beta,\mu$: constant stepsize $\eta$, momentum parametes $\beta,\mu$
@@ C.2 Additional details for synthetic experiments
[…]
@@ C.3 Additional details for decentralized training of Transformers
Transformer architecture. We consider the following decoder-only Transformer model (GPT): vocabulary size is 10208, context length is 64, embedding size is 128, number of attention heads is 4, number of attention layers is 2, the linear projection dimension within attention block is 512, and LayerNorm is applied after the 2nd attention block. The total number of parameters of this model is 3018240.
@@
DSGD-GClip $\alpha\in\{10^{-4},10^{-3},10^{-2},10^{-1},1,10,10^{2}\},\tau\in\{10^{-3},10^{-2},10^{-1},1,10,10^{2}\}$
DSGD-CClip $\alpha\in\{10^{-4},10^{-3},10^{-2},10^{-1},1,10,10^{2}\},\tau\in\{10^{-3},10^{-2},10^{-1},1,10,10^{2}\}$
@@
Hardware. We simulate the distributed training on one NVIDIA H100 GPU, using PyTorch 3.2 with CUDA 12. The total hyperparameter search and training procedure took around 100 GPU hours.